True or False? (1) The graph of y=3x+1 looks exactly like the graph ofy=2x , but shifted up 1 unit.
(2) The graph of any equation in the form y = mx + b passes through the point (0,0).
(3) The ordered pair (3,4) satisfies the equation 2y-3x=-6

Answer by tutorcecilia(2144)

(1) The graph of y=3x+1 looks exactly like the graph ofy=2x , but shifted up 1 unit.
ANSWER: False. y=3x+1 has a slope of 3/1 and the line is shifted 1 space to the left. y=2x has a slope of 2/1 and has a y-intercept of (0, 0).
.
.
(2) The graph of any equation in the form y = mx + b passes through the point (0,0). False. The answers will depending on the values of the slope (m) and the y-intercept (b).
.
.
(3) The ordered pair (3,4) satisfies the equation 2y-3x=-6.
Answer: Plug-in the values to test:
2(4)-3(3)=-6
8 - 9 does not equal -6
False.